# Generative Modelling of Structurally Constrained Graphs

### v1

@@ Abstract
Graph diffusion models have emerged as state-of-the-art techniques in graph generation, yet integrating domain knowledge into these models remains challenging. Domain knowledge is particularly important in real-world scenarios, where invalid generated graphs hinder deployment in practical applications. Unconstrained and conditioned graph generative models fail to guarantee such domain-specific structural properties. We present ConStruct, a novel framework that allows for hard-constraining graph diffusion models to incorporate specific properties, such as planarity or acyclicity. Our approach ensures that the sampled graphs remain within the domain of graphs that verify the specified property throughout the entire trajectory in both the forward and reverse processes. This is achieved by introducing a specific edge-absorbing noise model and a new projector operator. ConStruct demonstrates versatility across several structural and edge-deletion invariant constraints and achieves state-of-the-art performance for both synthetic benchmarks and attributed real-world datasets. For example, by leveraging planarity in digital pathology graph datasets, the proposed method outperforms existing baselines and enhances generated data validity by up to 71.1 percentage points.
@@ 1 Introduction
Figure 1: Constrained graph discrete diffusion framework. The forward process consists of an edge deletion process driven by the edge-absorbing noise model, while the node types may switch according to the marginal noise model. The projector operator ensures that sampled graphs remain within the constrained domain throughout the entire reverse process. In this case, the constrained domain consists exclusively of graphs with no cycles. We highlight in gray the components responsible for preserving the constraining property.
@@ 2 Related Work
By decomposing the graph generation task into multiple denoising steps, graph diffusion models have gained prominence due to their superior generative performance among methods that predict the full adjacency matrix at once (e.g., VAEs [39, 72, 76, 36], GANs [15, 41, 53], and normalizing flows [47, 52, 46, 50]). Diverse diffusion formulations have emerged to address different particularities of the graph setting, encompassing score-based approaches [59, 37, 83] and discrete diffusion [77, 26, 63]. They have also been used in intermediate steps for specific generation schemes, such as iterative local expansion [6]. Graph diffusion models are amenable to conditional graph generation techniques, which, despite enabling the guidance of the generation process towards graphs with desired properties, do not guarantee the satisfaction of such properties [77, 30]. Additionally, graph diffusion models can leverage formulations that are invariant to permutations, avoiding the sensitivity to node ordering that characterizes autoregressive models [85, 45, 13]. However, while the latter can straightforwardly ensure constraint satisfaction through validity checks at each iteration, the former still lag behind in addressing this challenge.
Previous approaches to close such gap can be categorized according to the assumed nature of graph state-spaces. In the continuous case, PRODIGY [70] exploits the projection of the sampled graph onto the constraint satisfying set at each reverse step. It can impose simple structural and molecular properties for which a closed-form projection can be derived. While this approach has shown success in other data modalities [11], applying a continuous state-space projection on the adjacency matrix and categorical node features can pervert their inherent discreteness and compromise the smoothness of the denoising process. Additionally, such an operator can not be derived to combinatorial constraints over the graph structure, as is often encountered in real-world scenarios (e.g., planarity or acyclicity).
@@
Then, in discrete state-spaces, EDGE [9] leverages a node-wise maximum degree hard constraint due to its degree guidance but it is limited to this particular property. Similarly, GraphARM [40], a graph autoregressive diffusion model, allows for constraint incorporation in the autoregressive manner. However, this method requires learning a node ordering, a task at least as complex as isomorphism testing. Therefore, to the best of our knowledge, ConStruct consists of the first graph constrained discrete diffusion framework covering a broad class of structural (potentially combinatorial) constraints.
@@ 3.1 Graph Diffusion Models
We first introduce the notation used in our model:
@@
We define a graph as $G=(X,E)$, where $X$ and $E$ denote the sets of attributed nodes and edges, respectively. We consider the node and edge features to be categorical and to lie in the spaces $\mathcal{X}$ and $\mathcal{E}$ of cardinalities $b$ and $c$, respectively. Thus, $x_{i}$ denotes the node attribute of node $i$ and $e_{ij}$ the edge attribute of the edge between nodes $i$ and $j$. With $\mathcal{H}^{k}=\{\mathbf{v}=(v_{1},\ldots,v_{k})\mid v_{i}\in\{0,1\},\sum_{i=1}^{k}v_{i}=1\}$, their corresponding one-hot encodings are then $\mathbf{x}_{i}\in\mathcal{H}^{b}$ and $\mathbf{e}_{ij}\in\mathcal{H}^{c+1}$, since we consider the absence of edge between two nodes as an edge type (“no edge” type). These are stacked in tensors $\mathbf{X}\in\{0,1\}^{n\times b}$ and $\mathbf{E}\in\{0,1\}^{n\times n\times(c+1)}$, respectively. So, equivalently to the set notation, we also have $G=(\mathbf{X},\mathbf{E})$.
@@
In some tasks, we are interested in only generating instances of a specific class of graphs that match the training distribution together with specific structural properties. This becomes particularly crucial, for instance, in scenarios where we possess domain knowledge but lack sufficient data for an unconstrained model to capture strict dependencies. It also applies to many real-world applications, where graphs that do not respect certain conditions are not feasible or have no physical meaning (e.g., drug design). Despite the remarkable expressivity of graph diffusion models, incorporating such constraints into their generative process remains a largely unsolved problem.
@@ 3.2 Constrained Graph Discrete Diffusion Models
We now introduce ConStruct, a framework that efficiently constrains graph diffusion models based on structural properties. Constraining graph generation implies guaranteeing that such target structural properties are not violated in the generated graphs. We build on graph discrete diffusion due to its intrinsic capability to effectively preserve fundamental structural properties (e.g., sparsity) of graphs throughout the generative process [77, 26].
@@
Many properties that are observed in real-world scenarios are edge-deletion invariant. For example, graph planarity is observed in road networks [81], chip design [7], biochemistry [71] or digital pathology [34]. In evolutionary biology [24] or epidemiology [68], we find graphs that must not have cycles. Also, maximum degree constraints are quite common in the design of contact networks [31, 1]. Finally, Definition 3.1 is extendable to continuous graph-level features through a binary decision (e.g., by thresholding continuous values into a boolean value).
@@ 3.3 Edge-deletion Aware Forward Process
with $\bar{\mathbf{Q}}^{t}_{X}=\mathbf{Q}^{1}_{X}\ldots\mathbf{Q}^{t}_{X}$ and $\bar{\mathbf{Q}}^{t}_{E}=\mathbf{Q}^{1}_{E}\ldots\mathbf{Q}^{t}_{E}$. Noising a graph amounts to sampling a graph from this distribution. For the nodes, we use the marginal noise model [77] due to its great empirical performance. Importantly, to preserve the constraining structural property throughout the forward process, and, consequently, throughout the training algorithm (see Algorithm 1 in Appendix), we propose the utilization of an edge-absorbing noise model [3]. This noise model forces each edge to either remain in the same state or to transition to an absorbing state (which we define to be the no-edge state) throughout the forward process. This edge noise model poses the forward as an edge deletion process, converging to a limit distribution that yields graphs without edges. Therefore, we obtain the following transition matrices:
@@
where $\alpha^{t}$ and $\alpha_{\operatorname{ABS}}^{t}$ transition from 1 to 0 with $t$ according to the popular cosine scheduling [58] and the mutual-information-based noise schedule ($\alpha^{t}=1-(T+t+1)^{-1}$[3], respectively. The vectors $\mathbf{1}_{b}\in\{1\}^{b}$ and $\mathbf{1}_{c}\in\{1\}^{c+1}$ are filled with ones, and $\mathbf{m}_{X}^{\prime}\in\Delta^{b}$ and $\mathbf{e}_{E}^{\prime}\in\{0,1\}^{c+1}$ are row vectors filled with the marginal node distribution and the one-hot encoding of the no-edge state, respectively.
@@ 3.4 Structurally-Constrained Reverse Process
The reverse process of the diffusion model is fully characterized by the distribution $p_{\theta}(G^{t-1}|G^{t})$. We detail how to build it from the predictions of a denoising graph neural network, (GNNθ), and the posterior term of the forward process in Section A.1. Importantly, the latter imposes the reverse process as an edge insertion process, yet does not necessarily ensure the target structural property. To handle that, we propose an intermediate procedure for each reverse step. Provided a noisy graph $G^{t}$ at timestep $t$, we do not accept directly $\hat{G}^{t-1}$, sampled from $p_{\theta}(G^{t-1}|G^{t})$, as the one step denoised graph. Instead, we iteratively add the newly added edges to $\hat{G}^{t-1}$ in a random order, discarding the ones that lead to the violation of the target property. Therefore, we only have $G^{t-1}=\hat{G}^{t-1}$ if none of the candidate edges breaks the target property. We refer to the operator that outputs $G^{t-1}$ provided $\hat{G}^{t-1}$ and $G^{t}$ by discarding the violating edges as the projector. Its implementation is illustrated in Figure 2 and described in Algorithm 3 in Appendix. Importantly, this procedure merely interferes with the sampling algorithm (refer to Algorithm 2 in Appendix) and ensures that the diffusion model training remains unaffected, preserving its efficiency.
@@
where $\mathcal{C}=\{G\in\mathcal{G}|P(G)=True,G\supset G^{t}\}$, with $\mathcal{G}$ the set of all unattributed graphs, then $G^{*}$ can be recovered by our projector, i.e., $G^{*}\in\mathcal{G}^{t-1}$.
The relationship between the projector and the candidate element $\hat{G}^{t-1}$ (the instance we aim to project onto a constrained set) and the specified target property $P$ (defining the constrained set) can be analogized to the conventional projection operator in continuous state spaces. However, while projection in continuous spaces is typically straightforward, this is not the case for discrete state spaces, where, for instance, there often lacks an inherent notion of order between different states. In particular, projecting into an arbitrary subclass of graphs is a complex general combinatorial problem to which there is no efficient solution. For example, finding the maximum planar subgraph of a given graph is NP-hard [10]. Therefore, the novelty of our method is introduced by considering an additional dependency on ${G}^{t}$: to make such problem efficiently approachable, we use the previous iterate, ${G}^{t}$, which we know by construction that verifies the target property, as a reference. This information is added into the optimization problem through the formulation of the set $C$. Importantly, this formulation is consistent with the designed noise for the diffusion model, as it complies with the reverse process as an edge insertion process (i.e., $G^{t}\subset G^{t-1}$). The complete version of this theorem and extensions for specific constraints can be found in Appendix B.
@@ 3.5 Implementation Improvements
Our reverse process consists solely of edge insertion steps, making it well-suited for the application of incremental algorithms. These algorithms efficiently check whether newly added edges preserve the target property by updating and checking smartly designed representations of the graph. This approach contrasts with full graph counterparts, leading to significant efficiency gains by reducing redundant computation. For instance, while the best full planar testing algorithm is $O(n)$ [29], its fastest known incremental test has amortized running time of $O(\alpha(q,n))$, where $q$ is the total number of operations (edge queries and insertions), and $\alpha$ denotes the inverse-Ackermann function [42] (often considered “almost constant” complexity). More details for different properties in Appendix C.
In practice, we observe the projector to incur a minor overhead in the runtime of the diffusion model sampling when compared to the model’s quadratic dependency on edge prediction for every timestep (see Section D.3).

### v2

@@ Abstract
Graph diffusion models have emerged as state-of-the-art techniques in graph generation; yet, integrating domain knowledge into these models remains challenging. Domain knowledge is particularly important in real-world scenarios, where invalid generated graphs hinder deployment in practical applications. Unconstrained and conditioned graph diffusion models fail to guarantee such domain-specific structural properties. We present ConStruct, a novel framework that enables graph diffusion models to incorporate hard constraints on specific properties, such as planarity or acyclicity. Our approach ensures that the sampled graphs remain within the domain of graphs that satisfy the specified property throughout the entire trajectory in both the forward and reverse processes. This is achieved by introducing an edge-absorbing noise model and a new projector operator. ConStruct demonstrates versatility across several structural and edge-deletion invariant constraints and achieves state-of-the-art performance for both synthetic benchmarks and attributed real-world datasets. For example, by incorporating planarity constraints in digital pathology graph datasets, the proposed method outperforms existing baselines, improving data validity by up to 71.1 percentage points.
@@ 1 Introduction
Figure 1: Constrained graph discrete diffusion framework. The forward process consists of an edge deletion process driven by the edge-absorbing noise model, while the node types may switch according to the marginal noise model. At sampling time, the projector operator ensures that sampled graphs remain within the constrained domain throughout the entire reverse process. In the illustrated example, the constrained domain consists exclusively of graphs with no cycles. We highlight in gray the components responsible for preserving the constraining property.
@@ 2 Related Work
By decomposing the graph generation task into multiple denoising steps, graph diffusion models have gained prominence due to their superior generative performance in the class of methods that predict the full adjacency matrix at once (e.g., VAEs [40, 74, 78, 37], GANs [16, 42, 54], and normalizing flows [48, 53, 47, 51]). Diverse diffusion formulations have emerged to address various challenges in the graph setting, encompassing score-based approaches [60, 38, 85] and discrete diffusion [79, 27, 64]. They have also been employed as intermediate steps in specific generative schemes, such as hierarchical generation through iterative local expansion [7].
The explicit incorporation of structural information (beyond local biases typical of GNNs) has been shown to be an important prior for enhancing the expressiveness of one-shot graph generative models. For example, in the GAN setting, SPECTRE [54] conditions on graph spectral properties to capture global structural characteristics and achieve improved generative performance. Graph diffusion models are similarly amenable to conditioning techniques [79, 31], which, despite enabling the guidance of the generation process towards graphs with desired properties, do not guarantee the satisfaction of such properties. In contrast, autoregressive models can ensure constraint satisfaction through validity checks at each iteration, effectively addressing this challenge. Although graph diffusion models can leverage formulations that are invariant to permutations, thus avoiding the sensitivity to node ordering that characterizes autoregressive approaches [87, 46, 14], they still lag behind in ensuring constraint satisfaction.
Previous graph diffusion approaches to address this challenge can be categorized according to the nature of the state spaces they assume. In the continuous case, aligned with successful outcomes in other data modalities [12], PRODIGY [72] offers efficient guidance for pre-trained models by relaxing adjacency matrices and categorical node features into continuous spaces, subsequently finding low-overhead projections onto the constraint-satisfying set at each reverse step. This approach can impose structural and molecular properties for which closed-form projections can be derived. However, it does not guarantee constraint satisfaction, facing a trade-off between performance and constraint satisfaction due to mismatched training and sampling distributions. This challenge arises from the continuous relaxation approach, which, while effective within the plug-and-play controllable diffusion framework, imposes an implicit ordering between states that can yield suboptimal graph representations when remapping to the inherently discrete graph space. Additionally, the proposed projection operators cannot be derived for some combinatorial constraints over the graph structure that are frequently encountered in real-world scenarios, such as planarity and acyclicity.
@@
Then, in discrete state-spaces, EDGE [10] leverages a node-wise maximum degree hard constraint due to its degree guidance but it is limited to this particular property. Similarly, GraphARM [41], a graph autoregressive diffusion model, allows for constraint incorporation in the autoregressive manner. However, this method requires learning a node ordering, a task that is at least as complex as isomorphism testing. Therefore, to the best of our knowledge, ConStruct consists of the first constrained graph discrete diffusion framework covering a broad class of structural (potentially combinatorial) constraints.
@@ 3.1 Graph Diffusion Models
We first introduce the mathematical notation adopted in the paper.
@@
We define a graph as $G=(X,E)$, where $X$ and $E$ denote the sets of attributed nodes and edges, respectively. We consider the node and edge features to be categorical and to lie in the spaces $\mathcal{X}$ and $\mathcal{E}$ of cardinalities $b$ and $c$, respectively. Thus, $x_{i}$ denotes the node attribute of node $i$ and $e_{ij}$ the edge attribute of the edge between nodes $i$ and $j$. With $\mathcal{H}^{k}=\{\mathbf{v}=(v_{1},\ldots,v_{k})\mid v_{i}\in\{0,1\},\sum_{i=1}^{k}v_{i}=1\}$, their corresponding one-hot encodings are then $\mathbf{x}_{i}\in\mathcal{H}^{b}$ and $\mathbf{e}_{ij}\in\mathcal{H}^{c+1}$, since we consider the absence of edge between two nodes as an edge type (“no edge” type). These are stacked in tensors $\mathbf{X}\in\{0,1\}^{n\times b}$ and $\mathbf{E}\in\{0,1\}^{n\times n\times(c+1)}$, respectively. So, equivalently to the set notation, we also have $G=(\mathbf{X},\mathbf{E})$. Additionally, we define the probability simplex, $\Delta^{k}=\{\left(\lambda_{0},\lambda_{1},\ldots,\lambda_{k-1}\right)\in\mathbb{R}^{k}\mid\lambda_{i}\geq 0\text{ for all }i,\text{ }\sum_{i=0}^{k-1}\lambda_{i}=1\}$.
@@
In some tasks, we are interested in generating instances of a specific class of graphs that conform to well-defined structural properties and align with the training distribution. Importantly, these structural properties do not fully define the underlying distribution; rather, the model must still learn this distribution within the specific class of graphs. This approach becomes particularly crucial in scenarios where we possess domain knowledge but lack sufficient data for an unconstrained model to capture strict dependencies, allowing us to reduce the task’s hypothesis space. This need also applies to many real-world applications, where generated graphs become irrelevant if they do not meet certain conditions, as they may be infeasible or lack physical meaning (e.g., in drug design). Despite the remarkable expressivity of graph diffusion models, incorporating such constraints into their generative process remains a largely unsolved problem.
@@ 3.2 Constrained Graph Discrete Diffusion Models
We now introduce ConStruct, a framework that efficiently constrains graph diffusion models based on structural properties. Constraining graph generation implies guaranteeing that such target structural properties are not violated in the generated graphs. We build on graph discrete diffusion due to its intrinsic capability to effectively preserve fundamental structural properties (e.g., sparsity) of graphs throughout the generative process [64, 79, 27].
@@
Many properties that are observed in real-world scenarios are edge-deletion invariant. For example, graph planarity is observed in road networks [83], chip design [8], biochemistry [73] or digital pathology [36]. In evolutionary biology [25] or epidemiology [70], we find graphs that must not have cycles. Additionally, if we consider the extensions of discrete diffusion to directed graphs (e.g., Asthana et al. [3]), there are several domains where graph acyclicity is critical: neural architecture search, bayesian network structure learning [88], or causal discovery [67]. Also, maximum degree constraints are quite common in the design of contact networks [32, 1]. Finally, it is worth noting that Definition 3.1 is extendable to continuous graph-level features through a binary decision (e.g., by thresholding continuous values into boolean values).
@@ 3.3 Edge-deletion Aware Forward Process
with $\bar{\mathbf{Q}}^{t}_{X}=\mathbf{Q}^{1}_{X}\ldots\mathbf{Q}^{t}_{X}$ and $\bar{\mathbf{Q}}^{t}_{E}=\mathbf{Q}^{1}_{E}\ldots\mathbf{Q}^{t}_{E}$. Noising a graph amounts to sampling a graph from this distribution. For the nodes, we use the marginal noise model [79] due to its great empirical performance. Importantly, to preserve the constraining structural property throughout the forward process, and, consequently, throughout the training algorithm (see Algorithm 1, in Section A.1), we propose the utilization of an edge-absorbing noise model [4]. This noise model forces each edge to either remain in the same state or to transition to an absorbing state (which we define to be the no-edge state) throughout the forward process. This edge noise model poses the forward as an edge deletion process, converging to a limit distribution that yields graphs without edges. Therefore, we obtain the following transition matrices:
@@
where $\alpha^{t}$ and $\alpha_{\operatorname{ABS}}^{t}$ transition from 1 to 0 with $t$ according to the popular cosine scheduling [59] and the mutual-information-based noise schedule ($\alpha^{t}=1-(T+t+1)^{-1}$[4], respectively. The vectors $\mathbf{1}_{b}\in\{1\}^{b}$ and $\mathbf{1}_{c}\in\{1\}^{c+1}$ are filled with ones, and $\mathbf{m}_{X}^{\prime}\in\Delta^{b}$ and $\mathbf{e}_{E}^{\prime}\in\mathcal{H}^{c+1}$ are row vectors filled with the marginal node distribution and the one-hot encoding of the no-edge state, respectively.
@@ 3.4 Structurally-Constrained Reverse Process
The reverse process of the diffusion model is fully characterized by the distribution $p_{\theta}(G^{t-1}|G^{t})$. We detail how to build it from the predictions of a denoising graph neural network, GNNθ, and the posterior term of the forward process in Section A.2. Importantly, the latter imposes the reverse process as an edge insertion process, yet does not necessarily ensure the target structural property. To handle that, we propose an intermediate procedure for each reverse step. Provided a noisy graph $G^{t}$ at timestep $t$, we do not accept directly $\hat{G}^{t-1}$, sampled from $p_{\theta}(G^{t-1}|G^{t})$, as the one step denoised graph. Instead, we iteratively insert the newly added edges to $\hat{G}^{t-1}$ in a random order, discarding the ones that lead to the violation of the target property. Therefore, we only have $G^{t-1}=\hat{G}^{t-1}$ if none of the candidate edges breaks the target property. We refer to the operator that outputs $G^{t-1}$ provided $\hat{G}^{t-1}$ and $G^{t}$ by discarding the violating edges as the projector. Its implementation is illustrated in Figure 2 and described in Algorithm 2, in Section A.3. Importantly, this procedure merely interferes with the sampling algorithm (refer to Algorithm 3, in Section A.3) and ensures that the diffusion model training remains unaffected, fully preserving its efficiency.
@@
where $\mathcal{C}=\{G\in\mathcal{G}|P(G)=True,G\supset G^{t}\}$ and $\mathcal{G}$ is the set of all unattributed graphs, then $G^{*}$ can be recovered by our projector, i.e., $G^{*}\in\mathcal{G}^{t-1}$.
The relationship between the projector, the candidate element $\hat{G}^{t-1}$ (the instance we aim to project onto a constrained set) and the specified target property $P$ (defining the constrained set) can be analogized to the conventional projection operator in continuous state spaces. However, while projection in continuous spaces is typically straightforward, this is not the case for discrete state spaces, where, for instance, there often lacks an inherent notion of order between different states. In particular, projecting into an arbitrary subclass of graphs is a complex general combinatorial problem to which there is no efficient solution. For example, finding the maximum planar subgraph of a given graph is NP-hard [11]. Therefore, the novelty of our method is introduced by considering an additional dependency on ${G}^{t}$: to make such problem efficiently approachable, we use the previous iterate, ${G}^{t}$, which we know by construction that verifies the target property, as a reference. This information is added into the optimization problem through the formulation of the set $C$. Importantly, this formulation is consistent with the designed noise for the diffusion model, as it complies with the reverse process as an edge insertion process (i.e., $G^{t}\subset G^{t-1}$). The complete version of this theorem and extensions for specific constraints can be found in Appendix B.
@@ 3.5 Implementation Improvements
Our reverse process consists solely of edge insertion steps, making it well-suited for the application of incremental algorithms. These algorithms efficiently check whether newly added edges preserve the target property by updating and checking smartly designed representations of the graph. This approach contrasts with full graph counterparts, leading to significant efficiency gains by reducing redundant computation. For instance, while the best full planar testing algorithm is $O(n)$ [30], its fastest known incremental test has amortized running time of $O(\alpha(\mathfrak{q},n))$, where $\mathfrak{q}$ is the total number of operations (edge queries and insertions), and $\alpha$ denotes the inverse-Ackermann function [43] (often considered “almost constant” complexity). More details for different properties in Appendix C.
At each reverse step, the denoising network makes predictions for all nodes and pairs of nodes. This results in $O(n^{2})$ predictions per step. Thus, the complexity of the sampling algorithm of the underlying discrete diffusion model is $O(n^{2}T)$. In addition, the complexity overhead imposed by the projector is $O(NV)$. Here, $V$ represents the complexity of the property satisfaction algorithm and $N$ is the total number of times this algorithm is applied throughout the reverse process. So, in total, we have $O(n^{2}T+NV)$. Our analysis in Appendix C shows that incremental property satisfaction algorithms have notably low complexity. For instance, in cases like acyclicity, lobster components, and maximum degree, we have $V=O(|E_{\textrm{added}}|)$. Since the projector adds one edge at a time, we have $V=O(1)$. Additionally, since the blocking edge hash table limits us to perform at most one property satisfaction check per newly proposed edge (either we have never tested it or it is already blocked), $N$ corresponds to the total number of different edges proposed by the diffusion model across the whole reverse process. A reasonable assumption is that the model proposes $N=O(|E|)$ edges throughout the reverse process, with $|E|$ referring to the number of edges of the clean graph. This is for example true if the model is well trained and predicts the correct graph. Most families of graphs are sparse, meaning that $O(|E|/n^{2})\to 0$ as $n\to\infty$. For example, planar and tree graphs can be shown to satisfy $|E|/n^{2}=O(1/n)$. Thus, we necessarily have $N\leq n^{2}$. For these reasons, we directly find $O(NV)\ll O(n^{2}T)$, highlighting the minimal overhead imposed by the projector compared to the discrete diffusion model. This explains the low runtime overhead observed for ConStruct, as detailed in Section D.3 (9% for graphs of the tested size). Therefore, we can conclude that asymptotically $O(n^{2}T+NV)=O(n^{2}T)$, i.e., the projector overhead becomes increasingly negligible relative to the diffusion algorithm itself as the graph size increases, highlighting the scalability of our method.